OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y <= x - 5 y <= -x - 4

OpenStudy (anonymous):

(1, 10) (-1, 10) (10, 1) (1, -10)

hartnn (hartnn):

plug in points in both the inequalities, if both gets satisfied, then the point lies in the solution set. example : point (2,5) x=2,y=5 x-5 = 2-5 =-3 y<=-3 ---> 5<=-3 is FALSE, hence 2,5 does not lie in the solution set.

hartnn (hartnn):

example 2: point (10,3) x=10,y=3 x-5 = 10-5 =5 y<=5 ---> 3<=5 is TRUE, hence 10,3 lies in the solution set. in this same way, check for all points.

OpenStudy (anonymous):

Im so confuse O.o

hartnn (hartnn):

lets check for 1st option : 1,10 y<=x-5 10<= 1-5 10<=-4 is FALSE , right ? so 1,10 does NOT lie in solution set. can you check for other choices ?

OpenStudy (anonymous):

I did this and got the second option am i right?

hartnn (hartnn):

no, try 3rd.

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

I have another question just like this one. Can you help? i still don't understand this

hartnn (hartnn):

do you have difficulty on how to plug in points ? or how to verify whether the point satisfies the inequality ?

OpenStudy (anonymous):

plugging in and everything that comes after that

OpenStudy (anonymous):

This is the question Which of the following points lie in the solution set to the following system of inequalities? y > -3x + 3 y > x + 2 (2, -5) (-2, 5) (2, 5) (-2, -5)

hartnn (hartnn):

Lets take first choice : (x,y)=(2,-5) ...x= 2, y=-5  can you try once to plug in the values? just replace y by -5 and x by 2, what u get ?